Number of colonies: Plate LB, no DNA LB/ amp, no DNA uncountable LB/ amp, DNA (The two plated labeled Bacteria DNA) 683 (70 ul) 291 (3011) (a) What quantity of DNA (in ng) was used in the transformation? (2 marks) (b) What fraction of the total transformation reaction was plated out? (2 marks) (c) How many transformants (colonies) were there in the total transformation reaction? (2 marks) (d) What was the transformation efficiency, expressed as transformants per ug of DNA added? (3 marks) If the highest efficiency obtainable by an optimal method is around 1 x 108 transformants per μg and a poor efficiency is around 1 x 10, how does your transformation efficiency compare? (1 mark)
2. You need 200 ml of 0.7% agarose to pour your gel. You are provided with agarose powder, a 5 x TBE solution, a 20,000X GelRed Nucleic Acid stain, and water. How much of each of these four components is required in your mixture to make the agarose gel? (8 marks) If you would like to run 20 ul of your PCR product in the above gel, how much 6X Gel Loading dye you need to prepare your sample? (2 marks) (Total 10 marks)
Answer #1

2.

  • Agarose = (percent)*(volume)

= 0.7*200

= 1.4g

  • The final concentration of TBE should be 1X for dissolving agarose.
    C1V1 = C2V2
    5*V1 = 1*200
    V1 = 200/5 = 40
    40mL of 5x TBE should be added
  • 200-40 = 160mL of water should be added to the mixture and the agarose should be dissolved.
  • Upon dissolution, once the agarose cools down a little, 10㎕ of 20000X GelRed Stain should be added. Based on the protocol provided by chembio, 5㎕ should be added to 100ml. Therefore, 10㎕ for 200mL should be added  

The concentration of loading dye should be 1X. Therefore, 4㎕ of 6X loading dye along with 20㎕ of the sample can be loaded.

1.

a) Usually, 1ng of DNA is used for transformation

For the other questions, the total reaction mixture is needed